OpenStudy (anonymous):

5/96+3sqrt2)-(2+2sqrt2) solve this please:)

OpenStudy (anonymous):

is that 5 /96 or 5(96) ?

OpenStudy (anonymous):

5 is over everything.. but its 5/(96+3sqrt2)-(2+2sqrt2)

OpenStudy (anonymous):

=5/(94+√2) =5/(94+√2) x (94-√2)/(94-√2) =5(94-√2) /(8836-2) =(470-5√2) /8834